I need to find the point of intersection between a circle and ellipse. Here are the equations:
x^2+y^2=625
(x+5)^2/9+(y-8)^2/100=1
I have no idea what to do and am about to break down!

Answer by lwsshak3(11628)

I need to find the point of intersection between a circle and ellipse. Here are the equations:
x^2+y^2=625
(x+5)^2/9+(y-8)^2/100=1
I have no idea what to do and am about to break down!
**
No wonder you are about to break down!
Punch these two equations into a graphing calculator and you will see that they do not intersect.
The ellipse is wholly inside the circle.